# Experimental Demonstration of a Quantum-Optimal Coronagraph Using Spatial Mode Sorters

### v1

@@ II.2 Measurement Model
with $b=1/N$ representing the relative brightness of the exoplanet. For our particular experimental setup we have integer constants $K=4$, $M=77^{2}$, and $N_{s}=1000$.
Figure 5: (a) Map of the log-likelihood $\mathcal{L}(\check{y}_{e};y_{e})$ over sub-diffraction exoplanet locations. The peak likelihood ridgeline (dark red regions) nearly coincides with perfect localization (black diagonal). However certain regions near the optical axis and outside the support of the truncated mode basis exhibit weakly-peaked likelihoods indicating greater MLE uncertainty. (b) Scatter plot of the MLEs obtained at each ground truth position. The colorbar indicates the frequency/density with which each estimate was made. Some outliers were removed for visual clarity.
Figure 6: (a) Mean empirical error of the exoplanet MLE as a function of the ground truth position (black curve). Green and blue shaded regions indicate star-planet separations where the localization error is below $\sigma/50$ and $\sigma/25$ respectively. (b) Statistical standard deviation of the MLE as a function of exoplanet position as compared to the Cramer-Rao lower bound (CRLB) for the measurement model of Equation 3. Uncertainty bars on the imprecision estimator were calculated using jackknife resampling Wu (1986). Red shaded region indicates viable bounds of the CRLB given uncertainty in the experimental model parameters. In particular we show the upper and lower bounds of the CRLB associated with the left and right sides of FWHM, respectively, for $\lambda_{0}$ in Figure 10 (Appendix C).
Figure 7: A comparison between theoretically predicted and measured intensity distributions for different sub-diffraction star-planet separations under a 1000:1 star-planet contrast. Theoretical images were simulated for an idealized version of our experimental system without background or modal crosstalk. In each image, the ground truth location of the exoplanet is coincides with the white crosshairs. The experimental measurements are shown after subtraction of the background. Asymmetries in the experimental intensity distributions for equidistant off-axis ±$\pm y_{e}$ exoplanet locations stem from modal cross-talk shown in Figure 4.
@@ III.1 Exoplanet Localization
In Figure 7, we compare simulated and experimental images of an artificial star-planet system captured with our coronagraph. For separations $|y_{e}|>0.1\sigma$, where the signal-to-noise ratio (SNR) of the exoplanet exceeds $\sim 1$, there are strong qualitative similarities between the simulated and experimental image intensity profiles, indicating that the exoplanet signal dominates over background noise. However, at separations below this threshold, where the vast majority exoplanet light is discarded with the fundamental mode, the noise in our experimental system becomes more prominent than the exoplanet signal. Additionally, we find that the asymmetry observed between images of the exoplanet positioned at equal distances above and below the optical axis (±$\pm y_{e}$) emerges due to modal cross-talk.
For each exoplanet location $y_{e}\in\mathcal{Y}$, we compiled repeated synthetic measurements $\mathbf{Y}^{(i)}(y_{e})$ for $i=1,\ldots,\ell$ (with $\ell=100$) with exposure time $T=0.1\text{s}$. To localize the exoplanet we employ a maximum likelihood estimator (MLE),
@@
Figure 5(a) shows a map of the log-likelihood $\mathcal{L}(y_{e};y_{e}^{\prime})=\log P\big{(}\mathbf{Y}(y_{e})|y_{e}^{\prime}\big{)}$ averaged over all experimental trials. The likelihood map exhibits a peak ridgeline (maximum likelihood) that corresponds almost exactly with the ground-truth position of the exoplanet, though exoplanet positions near the optical axis and outside of the modal region of support demonstrate greater estimator uncertainty (weakly-peaked likelihood). Figure 5b shows the MLEs obtained for all repeated measurements in the exoplanet translation scan. The estimator effectively localizes the exoplanet within the nominal region of support $|y_{e}|\leq.6\sigma$ of the truncated mode basis. Outside of this domain, the estimator experiences a bias due to the finite support of the truncated experimental mode set.
@@ III.2 Statistical Performance Analysis
To quantify the performance of our coronagraph, we analyze the statistical error and imprecision of the MLE over repeated experimental trials. For a given, ground truth exoplanet position, we denote the mean and variance of the MLE to be $\bar{y}_{e}\equiv\mathbb{E}_{\mathbf{Y}|y_{e}}[\check{y}_{e}]$ and $\sigma^{2}_{e}=\mathbb{V}_{{\mathbf{Y}|y_{e}}}[\check{y}_{e}]$ respectively. The unbiased empirical estimators of the MLE mean and variance are given by,
$\check{\bar{y}}_{e}(y_{e})$ $\equiv\frac{1}{\ell}\sum_{i=1}^{\ell}\check{y}_{e}^{(i)}(y_{e}),\,{\text{and}}$ (6a)
@@
Figure 6(a) shows the statistical error of the MLE over the complete domain of the exoplanet positional scan. In the subdomain $|y_{e}|\in[0.1,0.6]\sigma$, the absolute error is $|\check{\bar{y}}_{e}-y_{e}|<0.02\sigma$ with an estimator imprecision $\check{\sigma}_{y_{e}}\approx 0.01$. In the subdomain $|y_{e}|\in[.01,.1]\sigma$, the absolute error is $|\check{\bar{y}}_{e}-y_{e}|<0.03\sigma$ with imprecision $<\check{\sigma}_{y_{e}}\approx 0.1$. Figure 6(b) shows the empirical imprecision of the MLE as a function of the exoplanet position. We find that this empirical imprecision curve corresponds with the classical Cramer-Rao Lower Bound (CRLB) computed for our experimental measurement model. When the exoplanet is near the optical axis, most of the photons its photons couple to the fundamental mode and are discarded. Thus, the majority of detected photons are supplied by the background, giving rise to a central spike in the imprecision. Additionally, two secondary peaks in imprecision appear at outside the region of support for the truncated mode basis.
@@ IV Conclusion
In review, we have experimentally developed a quantum-optimal direct imaging coronagraph using spatial mode (de)multiplexing. We demonstrate high-accuracy localization of an artificial exoplanet at sub-diffraction distances from its host star under $1000:1$ star-planet brightness contrast. The performance of our experimental system at deeply sub-diffraction separations is fundamentally limited by a combination of detector dark noise, stray light, and modal cross-talk. At large star-planet separations our system performance is limited by the truncation of the mode basis, which may be improved by extending the number of MPLC phase planes. Overall, we believe this work further substantiates the potential value of mode-sorting solutions for astronomical imaging tasks.
Looking forward, even-order coronagraphs may be implemented using similar mode filtering principles presented here. Such a coronagraph would temper photon shot noise induced by a star of finite extent Belikov et al. (2021). Furthermore, extending functionality to broadband sources is necessary for analyzing the spectrum of the exoplanet. It is well-known that broadband sources introduce modal mismatch such that the cross-talk matrix becomes wavelength-dependent. For sub-diffraction exoplanets, the constraints on modal mismatch become more stringent Prasad (2020b); Grace (2022). Interestingly, preliminary numerical work suggests that the MPLC method may be used to simultaneously sort spectral and spatial modes Zhang et al. (2020), providing a common solution for exoplanet spectroscopy and/or broadband nulling of the star.
V Acknowledgements
@@ Appendix B Measurement Model
Here we expand on the single-shot measurement model of Equation 3. The post-nulling optical intensity distribution on the detector is modeled as,
$\mathbf{q}(\vec{r}_{0})=|\Psi\Omega C\Omega^{\dagger}\Psi^{\dagger}\psi_{0}(\vec{r}_{0})|^{2}$ (10)
which provides a matrix description of forward propagating the optical field through the MPLC, nulling the fundamental mode, back-propagating the field to an image plane, and measuring the intensity on a detector array. Here, $\psi_{0}(\vec{r}_{0})=\text{vec}[\bm{\psi}(\vec{r}-\vec{r}_{0})]$ is the vectorized form of the shifted PSF satisfying $\psi_{0}(\vec{r}_{0})^{\dagger}\psi_{0}(\vec{r}_{0})=1$. $\Psi$ is the truncated change-of-basis matrix for which the columns are the vectorized Fourier-Zernike modes $\psi_{k}=\text{vec}[\bm{\psi}_{k}(\vec{r})]$
@@
satisfying the orthonormality condition $\Psi^{\dagger}\Psi=I_{K}$ where $I_{K}$ is the identity on $\mathbb{C}^{K}$. The coronagraph matrix is given by,
$C=\begin{bmatrix}0&\mathbf{0}^{T}\\ \mathbf{0}&I_{K-1}\\ \end{bmatrix}$ (12)
which serves to null the fundamental mode (in the absence of cross-talk). While we treat the optical field as a scalar quantity (as the field is always is in a definite linear polarization state throughout our experiment), details on the experimental polarization manimuplation for path splitting are illustrated in Figure 8.
@@
Appendix C Calibration
This section addresses our procedures for characterizing all system-level parameters involved in the measurement model:
$\{\Omega,\lambda_{D},\lambda_{0},\mathbf{s}\}$
C.1 Cross-Talk Matrix $\Omega$
Figure 4 shows calibration measurements made for characterizing modal cross-talk through the MPLC. We estimate the cross-talk matrix from these calibration measurements by optimizing a least-squares objective function $\mathcal{F}(\Omega)$ under the constraint that $\Omega$ be unitary,
$\check{\Omega}=\mathop{\mathrm{argmin}}\limits_{\Omega\,\,:\,\,\Omega^{\dagger}\Omega=I_{K}}\underbrace{||M-(\Omega Z)\odot(\Omega Z)^{*}||_{F}^{2}}_{\mathcal{F}(\Omega)}$ (13)

### v3

@@ II.2 Measurement Model
where $b=1/N$ is the relative brightness of the exoplanet. Thus the synthetic measurement emulates the statistics of imaging two unbalanced incoherent point sources simultaneously through the coronagraph. In our experimental setup: $K=4$, $M=77^{2}$, and $N_{s}=1000$. Details regarding the empirical characterization of each parameter in the measurement model are provided in Appendix D.
II.3 Contrast Sensitivity Limit and Image Quality
The cross-talk matrix $\Omega$ of the mode sorter determines the performance of our coronagraph in two critical ways: First, energy leakage into and out of the fundamental mode sets a heuristic limit on the maximum accessible star-planet contrast. Second, interference with residual light in the fundamental mode causes distortions to the image of the exoplanet. Figure 5 shows the amplitudes and phase of each entry in the cross-talk matrix $\Omega$.
The maximum star-planet contrast accessible with our benchtop coronagraph is set by the leakage of the fundamental mode into higher-order modes. This leakage causes light from the on-axis star to corrupt otherwise pristine signal from the exoplanet. The throughput of the coronagraph is given by $P(\vec{r}_{0})=\mathbf{1}^{\intercal}\mathbf{q}(\vec{r}_{0})$, which represents the probability that a photon emitted by a point source at position $\vec{r}_{0}$ arrives at the detector after propagating through the system. The leakage out of the fundamental mode is given by the throughput of the on-axis star $P_{s}=P(\vec{0})$. Equivalently, the throughput for an exoplanet located at $\vec{r}_{e}$ is given by $P_{e}=P(\vec{r}_{e})$.
Under shot noise limited conditions, the SNR of the exoplanet signature is,
$SNR=\frac{\lambda_{e}P_{e}}{\sqrt{\lambda_{e}P_{e}+\lambda_{s}P_{s}}}$ (5)
where $\lambda_{e}$ and $\lambda_{s}$ are the mean photon fluxes of the exoplanet and the star respectively over the measurement period. Note that the SNR is maximal when there is no leakage out of the fundamental mode $P_{s}=0$. Additionally, we note that the shot noise contributed by the residual star light and by the exoplanet light are of similar magnitudes when the leakage equals the relative brightness $P_{s}\approx\frac{\lambda_{e}}{\lambda_{e}+\lambda_{s}}=b$. For this reason, the leakage sets a heuristic limit to the accessible levels of star-planet contrast. Using the experimental cross-talk matrix depicted in Figure 5, we find that the leakage for our benchtop system is $P_{s}=9.1975\times 10^{-4}$ which sets a contrast limit of approximately $b\gtrsim 10^{-3}$. In Figure 6, we show the measured throughput for our coronagraph and compare it with the modeled throughput predicted by the experimental cross-talk matrix.
The entries of the cross-talk matrix corresponding to imperfect isolation of the fundamental mode (i.e. first row/column of $\Omega$) also introduce distortions to the final image of the exoplanet. To see this, first consider an ideal cross-talk matrix of the form
$\Omega=\begin{bmatrix}1&\mathbf{0}^{\intercal}\\ \mathbf{0}&\tilde{\Omega}\end{bmatrix}$ (6)
where $\tilde{\Omega}\in\mathbb{C}^{(K-1)\times(K-1)}$ is a unitary sub-matrix describing the cross-talk between modes in the orthogonal complement space of the fundamental mode. Then, the details of $\tilde{\Omega}$ have no effect on the final image of the exoplanet since,
$\Omega C\Omega^{\dagger}=C$ (7)
such that Equation 2 reduces to $\mathbf{q}(\vec{r}_{0})=|\bm{\Psi}C\bm{\Psi}^{\dagger}\mathbf{u}(\vec{r}_{0})|^{2}$. This invariance of the nulling matrix $C$ under transformation by $\Omega$ reflects the fact that the PSF-adapted basis is not unique.
Conversely, if the mode sorter does not perfectly isolate the fundamental mode (i.e. $\Omega$ does not assume the form of Equation 6), then interference between the residual light in the fundamental mode and orthogonal modes cause distortions to the image of the exoplanet. The precise nature of these distortions depends on the specific amplitudes and phase of each off-diagonal entry in the cross-talk matrix given the target basis that the mode sorter is configured to (de)multiplex. In practice, Equation 7 must be approached with progressively higher fidelity to gain access to more extreme star-planet contrasts and lower image distortion. In Appendix E we perform a simulated analysis of the cross-talk matrix for the MPLC deployed in our experimental setup.
@@ III.1 Exoplanet Localization
In Figure 7, we compare theoretical, experimental, and simulated images of an artificial star-planet system captured with our coronagraph. For separations $|y_{e}|>0.1\sigma$ there are strong qualitative similarities between the theoretical and experimental image intensity profiles, indicating that the exoplanet signal dominates over the background noise. However, at separations below this threshold, where the vast majority of exoplanet photons are discarded with the fundamental mode, the inherent noise in our experimental system becomes more prominent than the exoplanet signal. Additionally, the asymmetry observed between images of the exoplanet positioned at equal distances above and below the optical axis (±$\pm y_{e}$) emerges due to the cross-talk between modes $\bm{\psi}_{1}$ and $\bm{\psi}_{2}$.
For each exoplanet location $y_{e}\in\mathcal{Y}$, we compiled repeated synthetic measurements $\{\mathbf{Y}^{(i)}(y_{e})\}$ for $i=1,\ldots,(\ell=100)$ with exposure time $T=100\text{ ms}$. To localize the exoplanet we employ a maximum likelihood estimator (MLE),
@@
We note that the MLE of the exoplanet position implicitly assumes a priori knowledge of the star-planet contrast $b$ via the distribution on $\mathbf{Y}$. In the idealized case where the cross-talk matrix takes the form of Equation 6 such that all light from the star is eliminated, knowledge of the contrast is not required because the measurement only contains contribution from the exoplanet $\mathbf{Y}\propto\mathbf{X}_{e}$. In practice, any true implementation of a mode sorter will likely suffer from some amount of cross-talk that prohibits complete extinction of the fundamental mode. Consequently, the star-planet contrast becomes a nuisance parameter for exoplanet localization and detection tasks. This reality applies to any coronagraphy system with leakage out of the fundamental mode, not specifically our design. A possible solution is to jointly estimate the exoplanet position and the contrast, though this extension lies beyond the scope of the current work.
Figure 8(a) shows a map of the log-likelihood $\mathcal{L}(y_{e};y_{e}^{\prime})=\log P\big{(}\mathbf{Y}(y_{e})|y_{e}^{\prime}\big{)}$ averaged over all experimental trials. The likelihood map exhibits a peak ridge line (maximum likelihood) that corresponds closely with the ground-truth position of the exoplanet. However, for exoplanets residing near the optical axis and outside the nominal region of support for the truncated mode basis, the likelihood exhibits weak curvature indicating greater estimator uncertainty. Figure 8(b) shows the MLEs obtained for all repeated measurements over the exoplanet translation scan, reinforcing salient features of the log-likelihood map. The MLE accurately localizes the exoplanet within the nominal region of support $|y_{e}|\leq.6\sigma$. Outside of this domain, the estimator experiences an increasing bias and uncertainty as photons are lost to modes beyond the support of the truncated mode set.
@@ III.2 Statistical Performance Analysis
Figure 8: (a) Map of the log-likelihood $\mathcal{L}(\check{y}_{e};y_{e})$ over sub-diffraction exoplanet locations. The peak likelihood ridge line (dark red regions) nearly coincides with perfect localization (black diagonal). However, certain regions near the optical axis and outside the support of the truncated mode basis exhibit weakly-peaked likelihoods indicating a greater variance in the MLE. (b) Scatter plot of the MLEs obtained from the complete measurement ensemble collected during the exoplanet position scan. The colorbar indicates the relative frequency of an estimate at each exoplanet position. Obvious outliers were removed for visual clarity.
To further quantify the performance of our coronagraph we analyze the statistical error and imprecision across the MLE realizations $\{\check{y}_{e}^{(i)}(y_{e}):y_{e}\in\mathcal{Y}\}$. Given a ground truth exoplanet position $y_{e}$, we denote the mean and variance of the MLE to be $\bar{y}_{e}(y_{e})\equiv\mathbb{E}_{\mathbf{Y}|y_{e}}[\check{y}_{e}]$ and $\sigma^{2}_{e}(y_{e})=\mathbb{V}_{{\mathbf{Y}|y_{e}}}[\check{y}_{e}]$ respectively. The unbiased empirical estimators of the MLE’s mean and variance are given by the statistical averages,
$\check{\bar{y}}_{e}(y_{e})$ $\equiv\frac{1}{\ell}\sum_{i=1}^{\ell}\check{y}_{e}^{(i)}(y_{e})$ (9a)
@@
The bias (expected error) of the MLE $\epsilon(y_{e})=\mathbb{E}_{\mathbf{Y}|y_{e}}[\check{y}_{e}]-y_{e}$ also has an associated estimator given by,
$\check{\epsilon}(y_{e})=\check{\bar{y}}_{e}(y_{e})-y_{e}$ (10)
Figure 9(a) shows the bias estimator (statistical error) of the MLE over the complete exoplanet position scan. In the subdomain $|y_{e}|\in[0.1,0.6]\sigma$, the absolute error is $|\check{\epsilon}(y_{e})|<0.02\sigma$ with an imprecision $\check{\sigma}_{e}(y_{e})\approx 0.01\sigma$, showcasing the localization accuracy achievable with our benchtop coronagraph at sub-Rayleigh star-planet separations. In the subdomain $|y_{e}|\in[0.01,0.1]\sigma$, where the experimental data is dominated by systematic error (primarily background and modal cross-talk), the absolute error is $|\check{\epsilon}(y_{e})|<0.03\sigma$ with a relatively large imprecision $\check{\sigma}_{e}(y_{e})\approx 0.1$ reflecting the fundamental uncertainty due to small signal in presence of significant measurement noise.
Figure 9: (a) Average error of the exoplanet MLE as a function of the ground truth position (black curve). Green and blue shaded regions demarcate where the absolute localization error is below $\sigma/50$ and $\sigma/25$ respectively. (b) Statistical standard deviation of the MLE as a function of exoplanet position as compared to the Cramér-Rao lower bound (CRLB) for the measurement model of Equation 3. Uncertainty bars on the imprecision estimator were calculated using jackknife resampling Wu (1986). The red shaded region indicates viable bounds of the CRLB given uncertainty in the experimental model parameters. In particular we show the upper and lower bounds of the CRLB associated with the left and right sides of FWHM, respectively, for $\lambda_{0}$ in Figure 12 (Appendix D).
The variance of the MLE is subject to the Cramér-Rao lower bound (CRLB) for a biased estimator,
$\sigma^{2}_{e}(y_{e})\geq\frac{(1+\epsilon^{\prime}(y_{e}))^{2}}{I(y_{e})}=\text{CRLB}(y_{e})$ (11)
where $\epsilon^{\prime}(y_{e})=\frac{d\epsilon}{dy_{e}}$ is the derivative of the bias and $I(y_{e})$ is the Fisher information of $y_{e}$ under the coronagraph measurement model. This bound sets the minimum imprecision for the MLE. Figure 9(b) shows the empirical imprecision $\check{\sigma}_{e}(y_{e})$ of the MLE as a function of the exoplanet position. We find that this imprecision curve corresponds well with the trends set by the CRLB of the MLE. As the exoplanet approaches the optical axis, a growing proportion of its photons couple to the fundamental mode and are discarded by the coronagraph. Thus, the majority of detected photons are supplied by the background, degrading the precision of the position estimate. Interestingly, when the exoplanet is nearly coincident with the optical axis, the MLE becomes a biased estimator (Appendix G) giving rise to a prominent dip in the imprecision. Additionally, we also observe apparent secondary lobes in the imprecision beyond the region of support due to mode truncation error. These lobes correspond to a decrease in the Fisher information of the exoplanet position evident in Figure 3(b).
@@ IV Conclusion
In summary, we present an experimental quantum-optimal direct imaging coronagraph design using spatial mode (de)multiplexing, which has inherent advantages for exoplanet exploration relative to indirect observation techniques. Furthermore, we demonstrate high-accuracy localization of an artificial exoplanet at sub-diffraction distances from its host star, which provides new capability to search and study exoplanets that may be inaccessible (i.e., in terms of inner working angle) to current state-of-the-art coronagraphs.
The performance of our experimental system at deeply sub-diffraction separations is presently limited by a combination of modal cross-talk, detector dark noise, and ambient background light. These experimental non-idealities can be mitigated comprehensively with higher-fidelity detectors, mode-sorter, and further optical design optimizations, enabling operation at deeper sub-diffraction separations and higher contrast levels. For large star-planet separations our system performance is limited by the truncation of the mode basis, which may be improved by extending the number of phase planes comprising the MPLC. Furthermore, recent techniques for in situ optimization of MPLC phase masks Rocha et al. (2025) could overcome model mismatch between our experimental system and its digital counterpart, enabling faithful numerical performance analyses. Additionally, while we employ a synthetic measurement construction in this work, follow-on studies involving two simultaneously radiating sources would provide more robust validation of our coronagraph design. Overall, we believe our work provides a compelling basis for further exploration and development of mode-sorting solutions for high-contrast astronomical imaging tasks.
Looking forward, even-order coronagraphs may be implemented using spatial mode filtering techniques similar to those presented here to accommodate the finite size of host star. Previous works have shown that shot noise induced by an extended host star may be progressively tempered by judiciously filtering/attenuating higher-order optical modes Belikov et al. (2021). Furthermore, extending functionality to broadband sources will be critical for analyzing the spectrum of the exoplanet, which embeds key planetary science information. Broadband sources introduce mode mismatch such that the cross-talk matrix becomes wavelength-dependent. For sub-diffraction exoplanets, the constraints on acceptable levels of mode mismatch become more stringent Prasad (2020b); Grace (2022) ostensibly limiting the efficacy of spatial mode sorting methods. However, recent numerical Zhang et al. (2020) and experimental Mounaix et al. (2020) work suggests that MPLCs may be used to simultaneously sort spectral and spatial modes, providing a common platform for exoplanet spectroscopy and/or nulling of a broadband star.
V Acknowledgments
@@
VI Disclosures
The authors declare no conflicts of interest.
VII Data Availability Statement
Our codebase for figure generation, coronagraph simulation, and measurement processing conducted in this work can be found in the project Github repository Deshler (a). Raw measurement data files are hosted on Zenodo and are free to download at Deshler (b).
@@ Appendix B Measurement Model
Here we expand on the single-shot measurement model of Equation 3. The post-nulling optical intensity distribution on the detector is given in Equation 2 to be,
$\mathbf{q}(\vec{r}_{0})=|\Psi\Omega C\Omega^{\dagger}\Psi^{\dagger}\mathbf{u}(\vec{r}_{0})|^{2}$
which provides a matrix description of forward propagating the optical field through the MPLC, nulling the fundamental mode, back-propagating the field to an image plane, and measuring the intensity on a detector array. Here, $\mathbf{u}(\vec{r}_{0})=\text{vec}[\bm{\psi}(\vec{r}-\vec{r}_{0})]$ is the vectorized form of the shifted PSF satisfying $\mathbf{u}(\vec{r}_{0})^{\dagger}\mathbf{u}(\vec{r}_{0})=1$. $\Psi$ is the truncated change-of-basis matrix for which the columns are the vectorized Fourier-Zernike modes $\psi_{k}=\text{vec}[\bm{\psi}_{k}(\vec{r})]$
@@
satisfying the orthonormality condition $\Psi^{\dagger}\Psi=I_{K}$ where $I_{K}$ is the identity matrix on $\mathbb{C}^{K}$. The coronagraph matrix is,
$C=\pm\begin{bmatrix}0&\mathbf{0}^{\intercal}\\ \mathbf{0}&I_{K-1}\\ \end{bmatrix}$ (16)
which serves to null the port assigned to the fundamental mode. The sign of coronagraph matrix depends on whether the system is transmissive (+)(+)( + ) at the sorting plane as in Figure 2(a) or reflective ()(-)( - ) at the sorting plane as in our experimental configuration depicted in Figure 2(b). While we treat the optical field as a scalar quantity (as the field is always is in a definite linear polarization state throughout our experiment), details on the experimental polarization manipulation for path splitting are illustrated in Figure 10.
@@
Appendix C Function Vectorization
Let $\Delta x,\Delta y$ be the horizontal and vertical pixel pitch dimensions for specific camera, and let $\{(x_{i},y_{i})\}_{i=1}^{M}$ be the set of points corresponding to the center of each pixel. The $\text{vec}(\cdot)$ operation takes a well-behaved function $f(x,y)$ over the plane and converts it to a discrete column vector $\mathbf{f}$ whose entries are given by,
[…]
Appendix D Calibration
This section addresses our procedures for characterizing all system-level parameters $\{\Omega,\lambda_{D},\lambda_{0},\mathbf{s}\}$ taken to be known quantities in the measurement model.
D.1 Cross-Talk Matrix $\Omega$
Figure 4 shows calibration measurements made at the sorting plane of the MPLC for characterizing modal cross-talk. For each position $y_{i}\in\mathcal{Y}$ of the point source, where $i=1,\ldots,(P=|\mathcal{Y}|)$, we measure the intensity in each mode channel $\mathbf{v}_{i}\in\mathbb{R}^{K}$. We estimate the cross-talk matrix from these calibration measurements by optimizing a least-squares objective function $\mathcal{F}(\Omega)$ under the constraint that $\Omega$ be unitary,
$\check{\Omega}=\mathop{\mathrm{argmin}}\limits_{\Omega\,\,:\,\,\Omega\in U(K)}\underbrace{||V-\tilde{V}(\Omega)||_{F}^{2}}_{\mathcal{F}(\Omega)}$ (18)